OpenStudy (anonymous):

A rocket is launched from atop a 105 foot cliff with an initial velocity of 156 ft/s . The height of the rocket above the ground at time t is given by h=-16t^2+156t+105. When will the rocket hit the ground after it is launched? Round to the nearest tenth. A) 4.9 s B) 9.8 s C) 0.6 s D) 10.4 s

OpenStudy (anonymous):

@♂

OpenStudy (anonymous):

When the rocket hits the ground, h=0. The equation then becomes\[-16t^2+156t+105=0\]Solve for t.

OpenStudy (anonymous):

so 4.9 ? @ospreytriple

OpenStudy (anonymous):

What are the two solutions you got for t by solving the quadratic?

OpenStudy (anonymous):

\[t = \frac{ 39 }{ 8 } - \frac{ \sqrt{1941} }{ 8 }\] \[t = \frac{ 39 }{ 8 } + \frac{ \sqrt{1941} }{ 8 }\]

OpenStudy (anonymous):

I used the quadratic formula and got t=-0.6 (discard, we can't have negative time) and t=10.38 which rounds to 10.4 waiting for osprey to finish before posting this so I'm not just spoonfeeding you

OpenStudy (anonymous):

i think i did it wrong

OpenStudy (anonymous):

thank you! @♂